OpenStudy (anonymous):

The equation y = 350x + 480 describes the amount of money in Blake’s savings account, where y is the amount of money in his account and x is the number of monthly deposits he makes. Eddie also has a savings account, and the balances in his account are recorded in the table. How much money did each account start with? Number of Monthly Deposits 1 2 4 7 Balance ($) 900 1375 2325 3750 Blake’s account started with $ . Eddie’s account started with $ .

OpenStudy (anonymous):

Eddie: 579? Blake: 357?

OpenStudy (texaschic101):

blake started with 480 (1,900) (2,1375) slope = (1375 - 900) / (2 - 1) slope = 475 y = mx + b 900 = 475(1) + b 900 = 475 + b 900 - 475 = b 425 = b y = 475x + 425 and Eddie started with 425

OpenStudy (anonymous):

Thank you @texaschic101 :) i will be sure to read it and learn how to do it. I hate it when people give direct answers because you dont learn anything. You are one of the few that give the explenation :)

OpenStudy (texaschic101):

basically, what they started with is your y intercept.... Blake started with 480 and added 350 each month Eddie started with 425 and added 475 each month

OpenStudy (anonymous):

:) thanks it helped alot